OpenStudy (anonymous):

y^n+y = sin x + cos 2x whats the general solution of this differential equation? thankssss

OpenStudy (anonymous):

(dn(x, y))/(dx) = (cos(x)-2 sin(2 x))/(y^n log(y))

OpenStudy (jamesj):

Do you mean the general solution to \[{d^n y \over dx^n } + y = \sin x + \cos 2x\] ?

OpenStudy (anonymous):

Ye

OpenStudy (jamesj):

Is this a problem you made up? Even writing down the general homogeneous/complimentary solution is a bit of pain. I think finding a general form for all n of the particular solution is going quite tricky.

OpenStudy (jamesj):

...going to be quite tricky

OpenStudy (jamesj):

Or perhaps not. Let's look at an Ansatz/trial solution for yp in the form yp = A sin x + B cos x + C sin 2x + D cos 2x It's not too hard to write down the nth derivative of this yp, with two cases: n even, n odd; and then solve for A, B, C, D

OpenStudy (jamesj):

The homogeneous part is likewise doable if we break it up into the cases where n is even and n is odd.